OpenStudy (anonymous):

The equation of a line is shown below: y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)? y = -x + 80 y = x - 80 y = 7x - 80 y = -7x + 80

OpenStudy (anonymous):

@mathmath333 @dan815

OpenStudy (anonymous):

@Conqueror @CausticSyndicalist

OpenStudy (anonymous):

if its parallel to that line what will its slope be?

OpenStudy (anonymous):

@heypeeps123

OpenStudy (anonymous):

y = -x - 7 This is in the form of y = mx + b, where 'm' is the slope. So what's the slope here? @heypeeps123

OpenStudy (anonymous):

 hey @heypeeps123 does it look like the drawing